A homeowner planning a kitchen remodeling can afford a $300 monthly payment. How much can the homeowner borrow for 2 years at 9%, compounded monthly, and still stay within the budget?
Answer #1

This requires application of PV of annuity formula,

PV =

1- (1+ г)-т Р т Р— Рeriodic Раутent т 3 гate рer period п 3D питber of periods

n = 2 * 12 months = 24 months, r = 9%/12 = 0.75% (per month)

PV = 300 * [\frac{1-(1+0.0075)^{-24}}{0.0075}]

PV = 300 * 21.8892

PV = $6,566.74